OpenStudy (anonymous):

What is the discharge of water moved by a river that has an area of 750m2 and is traveling 4 m/s?

OpenStudy (anonymous):

sumbody PLEASE help me :/

OpenStudy (kira_yamato):

A = 720 m^2 v = ds/dt = 4 m s^-1 dV/dt = 4*720 = 2880 m^3 s^-1

OpenStudy (anonymous):

that wasnt a option.... it was A. 3000 m3/s B. 750 m3/s C. 300 m3/s D. 4500 m3/s